OpenStudy (henrietepurina):

Esmeralda is graphing a polynomial function as a parabola. Before she begins graphing it, explain how to find the vertex. Make sure you include how to determine if it will be a maximum or minimum point. Use an example quadratic function to help you explain and provide its graph.

OpenStudy (henrietepurina):

@PaxPolaris @sleepyhead314 @TheSmartOne

OpenStudy (sleepyhead314):

umm two ways first way is to find the line of symmetry then plug that x value into the function and get the y value of the vertex point the other way is to complete the square and get it into the vertex form y = (x - h)^2 + k where (h, k) is the vertex point

OpenStudy (henrietepurina):

Would that be all, if I were to use that to answer?

OpenStudy (sleepyhead314):

no, I'm asking you to choose a method first for your answer

OpenStudy (henrietepurina):

Sure :)

OpenStudy (sleepyhead314):

If your equation is in standard form, then the formula for the axis of symmetry is: x = -b/2a from the general standard form equation y = ax2+bx + c

OpenStudy (henrietepurina):

Could I say that there are 2 ways?

OpenStudy (sleepyhead314):

mmm sure :)

OpenStudy (sleepyhead314):

and maximum/minimum of a vertex point depends on if the coefficient of x (the number attached to the x) is positive or negative if it is positive, then the vertex is a minimum if it is negative, then the vertex is a maximum

OpenStudy (henrietepurina):

Could my answer be: There are 2 ways to find the vertex. The first way is to find the line of symmetry then plug that x value into the function and get the y value of the vertex point. The other way is to complete the square and get it into the vertex form y = (x - h)^2 + k where (h, k) is the vertex point. If your equation is in standard form, then the formula for the axis of symmetry is: x = -b/2a from the general standard form equation y = ax2+bx + c. Also, and maximum/minimum of a vertex point depends on if the coefficient of x (the number attached to the x) is positive or negative: if it is positive, then the vertex is a minimum or if it is negative, then the vertex is a maximum.

OpenStudy (sleepyhead314):

I guess *shrugs* I don't know what your teacher wants... I'm just giving you what you need to answer the question whoops and I forgot a carrot in the standard form it's y = ax^2 + bx + c

OpenStudy (henrietepurina):

It's ok, would the answer above answer all of the questions asked?

OpenStudy (sleepyhead314):

all of it except for the example part

OpenStudy (henrietepurina):

What part is the example part?

OpenStudy (sleepyhead314):

Use an example quadratic function to help you explain and provide its graph.

OpenStudy (henrietepurina):

Ahh.. do we need it or do I need to remove it?

OpenStudy (sleepyhead314):

I can give you an example but you're going to have to try to do that explaining and graphing yourself :P

OpenStudy (sleepyhead314):

x^2 + 5x + 4 = y

OpenStudy (henrietepurina):

Hmm... what would I have to explain?

OpenStudy (sleepyhead314):

how to get the vertex from that and if it is a maximum or minimum

OpenStudy (henrietepurina):

Can you please help me do that?

OpenStudy (sleepyhead314):

using the information I gave can you find the line of symmetry?